OpenStudy (anonymous):

(4sqrt(2)+9)(sqrt(2)-1)

OpenStudy (anonymous):

\[\large (4 \sqrt{2}+9) \times (\sqrt{2}-1)\] \[\large F:4 \sqrt{2} \times \sqrt{2} \] \[\large O:4 \sqrt{2} \times -1 \] \[\large I:9 \times \sqrt{2} \] \[\large F:9 \times -1 \]

OpenStudy (anonymous):

FOIF ? that's sposed to be FOIL.... lol

OpenStudy (anonymous):

lol that is the right answer

OpenStudy (anonymous):

wrong